In a right triangle, one leg is 4 times the other. The hypotenuse squared is 153. what is the area of the triangle?
Answer by stanbon(75887)

In a right triangle, one leg is 4 times the other. The hypotenuse squared is 153. what is the area of the triangle?
---------------
Let one of the legs be "x"
The other is "4x"
EQUATION:
x^2 + (4x)^2 = 153
x^2 + 16x^2= 153
17x^2= 153
x^2=9
x=3
4x= 12
The area = (1/2)(base)(height)
area= (1/2)(3)(12)= 18 sq units.